### v4

# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

Abstract
This paper introduces a novel framework aimed at designing integrated waveforms for robust integrated sensing and communication (ISAC) systems. The system model consists of a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared-antenna deployment scenario. By leveraging Kullback-Leibler divergence (KLD) to holistically characterise both communication and sensing subsystems, three optimisation problems are formulated: (i) radar waveform KLD maximisation under communication constraints, (ii) communication waveform KLD maximisation subject to radar KLD requirements, and (iii) an integrated waveform KLD-based optimisation for ISAC that jointly balances both subsystems. The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby strengthening the trade-off between sensing and communication objectives and achieving higher overall performance than the individual radar- or communication-only techniques. Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem. Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming and identity-covariance radar baselines, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
Index Terms:
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, beamforming, optimisation, Kullback–Leibler divergence (KLD).
I Introduction
The rapid evolution of wireless communication networks has dramatically expanded the landscape of interconnected devices to include new applications such as autonomous vehicles and unmanned aerial vehicles (UAVs) that are heavily dependent on advanced sensing, pushing the boundaries of communication systems beyond their traditional limits [1, 2, 3, 4]. As we approach the sixth-generation (6G) era, network operators and service providers aim not only to further enhance their communication services but also to include advanced sensing-based applications such as detection, localisation, tracking, navigation, and environmental surveillance [5, 6, 7].
Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications. ISAC aims to synergistically merge sensing and communication functionalities hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilise resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems, which significantly enhance both communication capacity and sensing accuracy [16]. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics—communication performance through achievable rate, outage probability, and bit error rate (BER), and sensing performance through estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems. To this end, some unified performance measures have been introduced in the literature, such as the mutual information [18, 19, 20], and Kullback-Leibler divergence, also known as the relative entropy, which we recently introduced [21, 22, 23, 24].
I-A Literature review
Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25], where this approach minimises the CRB of angle estimation while maximising the signal-to-interference-plus-noise ratio (SINR) for communication. However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid designs to address hardware complexity issues [26]. While these approaches offer some practical advantages in terms of implementation, they fail to fully capture the nuanced trade-offs between sensing and communication performance in deeply integrated scenarios. Similarly, waveform optimisation has also been extensively studied, with recent work using faster-than-Nyquist approaches to enhance spectral efficiency while maintaining radar functionality as a constraint [27]. Recent studies have also investigated joint transmit and receive beamforming in full-duplex ISAC systems to manage interference and enhance overall performance; however, these approaches primarily emphasise interference suppression rather than exploitation, thereby neglecting potential performance gains achievable through direct integration of communication signals into radar sensing processes [28]. Despite these advances, existing work has not adequately addressed the need for a unified optimisation framework that simultaneously serves both subsystems’ requirements and balances their performance trade-offs effectively [29]. Furthermore, a recent comprehensive survey on interference management in ISAC systems highlights that most current strategies remain oriented toward interference mitigation or cancellation, reinforcing the perspective of communication signals as detrimental rather than beneficial [30]. Consequently, such designs not only overlook opportunities to leverage inherent signal interactions but also risk introducing excessive complexity, which is especially problematic when real-time processing is required in dynamic environments.
I-B Motivation and contribution
Despite these advances, current approaches often fall short of exploiting the full potential of the integrated system by assuming sensing and communication are competing subsystems rather than considering a synergic design. Most notably, there is a need for more comprehensive models that accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations, as shown in our paper, enhances radar detection capabilities, improving the system’s overall resource efficiency. This dual use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[31, 32].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario. Additionally, this work investigates the trade-off between trade-offs between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [22]. Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model. To this end, the key contributions of this work are:
  • Two Novel Separate Waveform Optimisation Designs: we present radar and communication waveform optimisation using the unified KLD measure to represent the achievable performance for both subsystems on one scale. Both formulated problems aim at maximising the KLD measure of a certain subsystem subject to the minimum KLD of the other subsystem and power constraints. The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex nature, enabling balanced performance optimisation across both subsystems within system constraints.
  • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together. This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, leading to enhanced performance in terms of radar detection accuracy and communication reliability compared with the separate waveform methods. By capturing and exploiting the synergy between sensing and data transmission, the unified technique achieves higher overall performance in shared-antenna ISAC deployments.
  • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection, which offers new insights into the underlying synergy between radar and communication waveforms and quantifies how their joint optimisation can enhance overall ISAC capabilities.
  • Comprehensive Complexity Assessment: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility. This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline. The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, with modest radar performance gains. Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency. The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering even higher overall performance in terms of sensing and communication signals. The numerical results highlight how leveraging communication signals for radar detection significantly enhances overall system performance, especially in scenarios with constrained dedicated radar resources. These findings emphasise the practical significance of KLD-based designs for shared-antenna ISAC systems, provide important adaptability and insights for balancing sensing and communication requirements in future wireless networks.
I-C Paper organisation
The paper is structured as follows. Section II presents the system model. Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the radar waveform optimisation is investigated using the KLD. Section VI, the communication waveform optimisation is investigated using the KLD. Section VII shows the complexity analysis for both waveform optimisation techniques, Section VIII displays the numerical results, and finally, Section IX concludes the work.
$Notation$: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors. Superscripts ${(\cdot)}^{*}$, ${(\cdot)}^{T}$, and ${(\cdot)}^{H}$, denote the conjugate, transpose, and Hermitian transpose, respectively. Subscripts ${(\cdot)}_{\mathrm{c}}$ and ${(\cdot)}_{\mathrm{r}}$ relate to the communication, and radar subsystems, respectively.
II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},$ (1)
where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
II-A Communication System
At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$. The received signal at the $k$-th UE can be formulated as
$y_{\mathrm{c},k,l}=\overbrace{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}}^{\text{Desired $k$-th UE signal}}+\overbrace{\omega_{\mathrm{MN},k}}^{\begin{subarray}{c}\text{Inter-user}\\ \text{interference}\end{subarray}}+\overbrace{\eta_{k}}^{\begin{subarray}{c}\text{Radar interference}\\ \text{and noise}\end{subarray}},$ (2)
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. ZF designs which is typically used to precode the communication data symbols at the BS. This precoding scheme will serve as a benchmark for our designs. The ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (3)
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
II-B Radar system
The radar system is designed for high adaptability, allowing for real-time adjustments in detection capabilities across $L$ frame snapshots. $T$ represents the radar beams emitted during a specific detection frame. Using MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [33, 34], where $T$ signifies the upper limit of detectable targets in a single frame.
The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification. Its flexibility in varying emitted beams allows for the simultaneous detection of more targets, to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems. Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [35, 36]  [^1] The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [37].. The total radar return signal at the $l$-th snapshot is shown as follows,
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vspace{-0.05 in}$ (4)
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix that captures the complete BS-Target-BS path characteristics for the $t$-th target, inherently incorporating propagation effects, target cross-section, and reflection coefficients. The target response matrix follows a Rayleigh distribution to model the composite effects of multiple scattering centres inherent in extended targets, thereby characterising the statistical nature of target scattering while preserving the direct path essential for target detection [38, 39]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot. To process the radar returns from all beams, a bank of matched filters is typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Assuming $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$ using signal space and time domain, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target [39, 34]. The radar precoding matrix $\mathbf{W}_{\mathrm{r},l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS after match-filtering, under the binary hypothesis testing problem $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ represents the target presence whereas $\mathcal{H}_{0}$ denotes the target absence, is denoted by $\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}$ and can be mathematically represented as follows,
$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}\>q+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}\>q+\mathbf{n}_{t,l},$ (5)
where $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot at the output of the matched filter.
III The communication KLD: ZF and Arbitrary precoding matrix
In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver. The analysis is carried out for the normalised precoders introduced in Section II-A. To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right).$ (6)
III-A ZF beamforming
With ZF precoding, the received signal at the $k$-th user is,
$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l},$ (7)
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix. Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (8)
where $\mathbf{R}_{w,t}\!\!=\!\!\frac{1}{L}\!\sum_{l=1}^{L}\!\!\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}^{H}$. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
$\boldsymbol{\mu}_{k,j}$ $=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]s_{\mathrm{c},k,l}^{(j)}$
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (9)
$\ \Sigma_{n}$ $=\Sigma_{m}=\sigma_{\eta}^{2},$ (10)
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (9), and (10) in (6), and considering all possible pairs of dissimilar symbols [22, Corollary 1], then the KLD becomes,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}}{M(M-1)\sigma_{\eta}^{2}\ln 2}\mathbb{E}\left[\frac{1}{\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}}\right],$ (11)
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is a constant that depends on the constellation. The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
[…]
Substituting this result into (11), we get the final $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N-K)/M(M-1)\sigma_{\eta}^{2}\ln 2.$ (13)
III-B Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem. Considering the received signal representation in (2), the conditional mean vectors and variance are,
$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\$ (14)
$\operatorname{Var}(y_{\mathrm{c},k,l}^{(j)})$ $=\mathbb{E}[|y_{\mathrm{c},k,l}^{(j)}-\boldsymbol{\mu}_{k,n}|^{2}]$
$=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\quad j\in\{n,m\},$ (15)
Note that, the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},$ (16)
by substituting the evaluated mean vectors and covariance matrices in (14), and (16) into the KLD expression given in (6), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, which can be represented as follows,
$\!\!\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}\!=\!\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
×$\!\!\times\!\Bigg{(}\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg{)}^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg{)}\!.\!\!$ (17)
Thereafter, by considering all possible pairs of dissimilar symbols, then the KLD expression using $\lambda$ is simplified as,
$\begin{split}\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (18)
It should be noted that the second term in this equation, represents a form of the SINR, where the interference is the sum of IUI and radar interference.
IV Radar system analysis
IV-A Composite detection and response matrix estimation
The main objective of typical radar systems is to detect the existence of targets and estimate the matrix response for existing targets. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{q}).$ (19)
This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining ^$\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if presents. It can be observed from (5) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t,l}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
×$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, which is derived as,
$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (21)
As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem, we obtain,
$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1}).$ (22)
Since the PDF of $\mathbf{y}_{\mathrm{r},t,l}$ is multivariate Gaussian, the maximisation leads to the least squares solution given by,
$\hat{\mathbf{H}}_{t}\!$ $=\!\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}\!-\!\mathbf{H}_{t}\mathbf{x}_{t,l}\right\|^{2}\!=\!\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t,l}^{H}\left(\mathbf{x}_{t,l}\mathbf{x}_{t,l}^{H}\right)^{-1}.$ (23)
This least squares solution provides an unbiased estimate of the target‐response matrix $\mathbf{H}_{t}$, with accuracy directly tied to the detection probability $P_{D}$. When the radar correctly detects a target under $\mathcal{H}_{1}$, the maximum likelihood solution $\hat{\mathbf{H}}_{t}$ is used to estimate $\mathbf{H}_{t}$. In contrast, a miss detection (i.e., deciding $\mathcal{H}_{0}$ under $\mathcal{H}_{1}$) forces $\hat{\mathbf{H}}_{t}=\mathbf{0}$, resulting in significant errors. The overall unconditional MSE is defined as,
$\mathrm{MSE}$ $=p(H_{1})\Bigl{[}(1-P_{D})\,\mathrm{MSE}_{0\_1}+P_{D}\,\mathrm{MSE}_{1\_1}\Bigr{]}$
$+\bigl{[}1-p(H_{1})\bigr{]}\Bigl{[}P_{\mathrm{FA}}\,\mathrm{MSE}_{1\_0}\Bigr{]},$ (24)
[…]
Substituting the PDF in (20) into the GLRT, we obtain,
$\Lambda(\mathbf{y}_{\mathrm{r},t,l})=$
$\frac{\!\sigma_{n}^{2N}\!\exp\!\left(\!-(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})\!\right)}{\!\det(\hat{\mathbf{R}}_{1,t})\exp\!\left(\!-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\!\right)}\!\!\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\!\tau_{0}$ (25)
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold. Taking the logarithm of the likelihood ratio and rearranging terms, the final GLRT form is given by,
$\dot{\Lambda}(\mathbf{y}_{\mathrm{r},t,l})=\mathbf{y}_{\mathrm{r},t,l}^{H}\left(\mathbf{I}_{N}+\hat{\mathbf{R}}_{1,t}^{-1}\right)\mathbf{y}_{\mathrm{r},t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{1}$ (26)
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. Here, the constant terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ have been absorbed into the threshold. This expression decides the most probable hypothesis, i.e., $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$, by incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. Since the received signals from different snapshots are conditionally independent, (26) can be updated to adapt to $L$ snapshot scenario by taking the average, i.e., $\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}$, with $\tau_{2}$ is the detection threshold considering all snapshots and can be designed according to Neyman-Pearson lemma.
IV-B The KLD of radar system
[…]
Using this property, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (20), and (6) as follows,
$\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=$ $\frac{1}{\ln{2}}\Big{(}\ln\left(\mathrm{det}(\mathbf{R}_{2,t})\right)+\mathrm{Tr}\left(\mathbf{R}_{2,t}^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)$
$-N\>\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\Big{)},$ (27)
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$. The average KLD of the whole ISAC system can be formulated as follows,
$\mathrm{KLD}_{\mathrm{ISAC}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}+\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}.$ (28)
V Radar Waveform Optimisation
In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (29b)
${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (29c)
$\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{r}},$ (29d)
$P_{\mathrm{c}}=P_{\mathrm{T}}-P_{\mathrm{r}},$ (29e)
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (27). After substituting (27) in (29a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively. The constraint on radar power (29d) is explicit because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (29c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
$\mathcal{P}_{2}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}$ $\!\!\!\!\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\ln{2}}\Bigg{(}\!\ln\!\left(\mathrm{det}(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N})\!\right)$ (30)
$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\!\sigma_{n}^{2}\!\right)\!-\!N\!\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\!\Bigg{)}$
s.t. (29b), (29c), (29d), and (29e)
To solve the optimisation problem in (30), we utilise the Projected Gradient method with a penalty function [40]. This method is chosen for its suitability to the non-linear, non-convex optimisation problem [41]. This method uses a penalty function for KLD constraints and handles radar power via projection, balancing the solution quality and computational efficiency for real-time ISAC applications. Its scalability makes it suitable for radar waveform optimisation in ISAC. The optimisation variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor. We define the objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ and a penalty function $p(\mathbf{\dot{W}}_{\mathrm{r}})$ to handle the constraints in (29b) and (29c),
$f(\mathbf{\dot{W}}_{\mathrm{r}})=\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}},$ (31)
$\begin{split}p(\mathbf{\dot{W}}_{\mathrm{r}})=&\sum_{t=1}^{T}\left(\max(0,\underline{A}_{t}-\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}})^{2}\right)\\ &+\sum_{k=1}^{K}\left(\max(0,\underline{B}_{k}-\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}}})^{2}\right).\end{split}$ (32)
The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix I.A. The final equation is given by,
$\begin{split}[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=&\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}\Big{(}\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ &-\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big{)}_{i}.\end{split}$ (33)
The gradient of the penalty function $\nabla p(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$ is also a gradient tensor, computed as,
$\begin{split}[\nabla p(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=&-2(\underline{A}_{t}-\mathrm{KLD}_{\mathrm{r},t})[\nabla\mathrm{KLD}_{\mathrm{r},t}]_{i,l}\cdot\mathbb{I}_{1}\\ &-\sum_{k=1}^{K}2(\underline{B}_{k}-\mathrm{KLD}_{\mathrm{c},k})[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}\cdot\mathbb{I}_{2},\end{split}$ (34)
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A. The final forms are given by,
$[\nabla\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}]_{i,l}=$ $\frac{2\>d_{\mathrm{r},t}^{\zeta}}{L\ln{2}}\Big{(}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$
$-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big{)}_{i},$ (35)
$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}\!=\!$ $-\!\frac{2d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N\!\!-\!\!K)\sigma_{h}^{2}}{M(M\!\!-\!\!1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\!\mathrm{r}}]_{i,l,t}.$ (36)
The update rule for the projected gradient method with penalty ensures that each iteration moves the solution towards optimality while maintaining feasibility, and can be shown as follows,
$\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}+\alpha_{n}\mathbf{G}^{(n)}\right),$ (37)
where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, balancing the objective and constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (29d) is always satisfied. The projection mechanism is shown as,
$\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}})=\begin{cases}\mathbf{\dot{W}}_{\mathrm{r}}\sqrt{\frac{P_{\mathrm{r}}L}{\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}}},&\text{if }\frac{1}{L}\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}>P_{\mathrm{r}}\\ \mathbf{\dot{W}}_{\mathrm{r}},&\text{otherwise}\end{cases}$ (38)
Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimisation
0:  Initial point $\mathbf{\dot{W}}_{\mathrm{r}}^{(0)}$, constants $\alpha_{0}$, $\rho_{0}$, $\beta$, $\gamma$, $c$, $P_{\mathrm{r}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\varepsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{\dot{W}}_{\mathrm{r}}^{*}$
1:  Initialise $n=0$
2:  while $n<\mathrm{max\_iter}$ do
3:     Compute gradients $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ and $\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$
4:     Compute total gradient $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$
5:     Perform backtracking line search to find $\alpha_{n}$
6:     Update $\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}=\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}+\alpha_{n}\mathbf{G}^{(n)})$
7:     Compute $f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})$ and $p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})$
8:     Update $\rho_{n+1}$ according to the penalty update rule
9:     if $\|\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\|_{F}<\varepsilon$ then
10:        break
11:     end if
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{\dot{W}}_{\mathrm{r}}^{*}=\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}$
The step size $\alpha_{n}$ is determined using a backtracking line search to ensure convergence. We find the smallest non-negative integer $m$ such that,
$\begin{split}&f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\geq\\ &f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})+c\alpha_{n}\|\mathbf{G}^{(n)}\|_{F}^{2},\end{split}$ (39)
where $c\in(0,1)$. This condition, known as the Armijo condition, ensures a sufficient decrease in the penalised objective function [42]. The step size is set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, with $\beta\in(0,1)$ as a reduction factor. The penalty parameter $\rho_{n}$ is updated as follows,
$\rho_{n+1}=\begin{cases}\gamma\rho_{n},&\text{if }p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})>0\\ \rho_{n},&\text{otherwise}\end{cases}$ (40)
where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied. The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
$\begin{split}\!\!\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{(n+1)})\!\!\leq&\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{\!(n)})\!+\!c\alpha_{n}\!\langle\nabla\!\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{\!(n)}),\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}\!\!-\!\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\!\!\!\end{split}$ (41)
where $\phi(\mathbf{\dot{W}}_{\mathrm{\!r}})\!=\!\!f(\mathbf{\dot{W}}_{\mathrm{\!r}})\!-\!\!\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{\!r}})$ is the penalised objective function. The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation, allowing the algorithm to balance optimisation and constraint satisfaction. If constraints are consistently violated, the increasing penalty will emphasise constraint adherence.
VI Communication optimisation
The communication subsystem optimisation uses the gradient-assisted IPM [43], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures. Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions. The communication waveform is optimised while using the conventional identity covariance (CIC) matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be expressed as follows,
$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (42a)
s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (42b)
${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (42c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (42d)
$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.$ (42e)
Note, that the radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow. To solve this optimisation problem, we must derive the gradients of the objective function and the constraints. Let’s define the objective function as follows,
$f(\mathbf{W}_{\mathrm{c}})=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}.$ (43)
To compute the gradient of the objective function, we need to derive the gradients of each component of $\mathbf{W}_{\mathrm{c}}$. Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$. Using the chain rule and properties of matrix derivatives, the gradients can be expressed as follows,
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\quad\quad\quad\quad\forall j$ (44)
$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ =$=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (45)
where $C=\frac{\lambda}{2M(M-1)\ln 2}$, and $b_{k}=d_{\mathrm{c},k}^{\zeta}\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$. Finally, we can express the gradient of the objective function as,
$\begin{split}\nabla_{\mathbf{w}_{\mathrm{c},k}}f(\mathbf{W}_{\mathrm{c}})=\frac{d_{\mathrm{c},k}^{\zeta}C}{K}\sum_{k=1}^{K}\Bigg{(}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}}\\ -\sum_{j\neq k}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}}\Bigg{)}.\end{split}$ (46)
Next, we need to consider the constraints of our optimisation problem. The power constraint is given by,
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vspace{-0.05 in}$ (47)
The gradient of this constraint is derived as follows,
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vspace{-0.05 in}$ (48)
The gradient of the radar KLD constraint with respect to $\mathbf{W}_{\mathrm{c}}$ is derived as follows,
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vspace{-0.03 in}$ (49)
Algorithm 2 Interior Point Method for Communication Optimisation
0:  Initial point $\mathbf{W}_{\mathrm{c}}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{W}_{\mathrm{c}}^{*}$
1:  Initialise $n=0$, $\mu=\mu_{0}$
2:  while $n<\mathrm{max\_iter}$ do
3:     Compute $f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $\nabla f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$
4:     Compute KKT conditions
5:     if KKT conditions satisfied within $\epsilon$ then
6:        break
7:     end if
8:     Solve Newton system for search direction $\Delta\mathbf{W}_{\mathrm{c}}$
9:     Perform line search to find step size $\alpha$
10:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}=\mathbf{W}_{\mathrm{c}}^{(n)}+\alpha\Delta\mathbf{W}_{\mathrm{c}}$
11:     Update $\mu=\gamma\mu$
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
Solving the optimisation problem defined in equations (42a)-(42d) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
VII Integrated Waveform Optimisation for ISAC
Finally, we introduce this unified approach that maximises the whole ISAC system KLD, which ensures improved radar detection and communication reliability while satisfying minimum subsystem KLD constraints and adhering to a predefined power budget. The optimisation problem is expressed as,
$\mathcal{P}_{1}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\quad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (50a)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t\in T,$ (50b)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k\in K,$ (50c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (50d)
$P_{\mathrm{T}}=P_{\mathrm{r}}+P_{\mathrm{c}}.$ (50e)
The total power constraint in (50d) limits the combined radar and communication power to $P_{\mathrm{T}}$. To solve this non-convex problem, the ADMM is employed. This approach unifies the projected gradient method used for radar waveform optimisation and the IPM technique used for communication waveform optimisation, efficiently handling the coupling between subsystems while decomposing the complex problem into more manageable subproblems. This involves introducing auxiliary variables $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ such that $\mathbf{Z}_{\mathrm{r}}=\mathbf{W}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}=\mathbf{W}_{\mathrm{c}}$. The augmented Lagrangian for the problem is given as,
$\mathcal{L}_{\rho}=$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}+\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}$
$+\langle\mathbf{U}_{\mathrm{r}},\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\rangle+\langle\mathbf{U}_{\mathrm{c}},\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\rangle$
$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),$ (51)
where $\mathbf{U}_{\mathrm{r}}$ and $\mathbf{U}_{\mathrm{c}}$ are the dual variables (Lagrange multipliers), and $\rho>0$ is the ADMM penalty parameter. The ADMM algorithm proceeds iteratively, updating the primal variables $(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$, auxiliary variables $(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$, and dual variables $(\mathbf{U}_{\mathrm{r}},\mathbf{U}_{\mathrm{c}})$. The updates are given as follows,
$\mathbf{W}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{r}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{r}}^{(n)},$ (52)
$\mathbf{W}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{c}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{c}}^{(n)},$ (53)
$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\quad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (54)
$\mathbf{U}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{r}}^{(n)}+\rho(\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}),$ (55)
$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).$ (56)
The auxiliary variables are updated by solving the following constrained subproblem,
$\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\quad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (57)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t,$ (58)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k,$ (59)
$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},$ (60)
where $p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$ is a penalty term that ensures constraints are satisfied. Since the subproblem (57) may not have a closed-form solution, it is solved approximately using projected gradient descent. The auxiliary variables are iteratively updated as: $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathbf{Z}_{\mathrm{r}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{r}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{r}}}p\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathbf{Z}_{\mathrm{c}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{c}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{c}}}p\right)$, where $\rho_{n}$ is the penalty parameter at iteration $n$, and $\alpha_{n}$ is the step size determined by a backtracking line search. The gradients of $f$ with respect to $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ are computed as,
$\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (61)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).$ (62)
The closed-form expressions for $\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ and $\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ are given by,
$\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (63)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ $=\frac{2\lambda d_{\mathrm{c},k}^{-\zeta}}{M(M-1)\ln 2}\Bigg{[}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},k}}{\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}}$
$\quad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg{]}.$ (64)
After each gradient-based update, the auxiliary variables are projected onto the feasible set defined by the power constraint as, $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right)$, where $\mathcal{P}(\cdot)$ denotes the projection operator ensuring that (60) is satisfied. This iterative procedure continues until convergence criteria, such as primal and dual residual thresholds, are met.
Algorithm 3 ADMM for Unified ISAC Waveform optimisation
0:  Initial variables $\mathbf{W}_{\mathrm{r}}^{(0)}$, $\mathbf{W}_{\mathrm{c}}^{(0)}$, $\mathbf{Z}_{\mathrm{r}}^{(0)}$, $\mathbf{Z}_{\mathrm{c}}^{(0)}$, $\mathbf{U}_{\mathrm{r}}^{(0)}$, $\mathbf{U}_{\mathrm{c}}^{(0)}$, parameters $\rho$, $\alpha_{0}$, $\rho_{0}$, tolerance $\varepsilon$, max iterations $\mathrm{max\_iter}$
0:  Optimized waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
1:  Initialize $n=0$
2:  while $n<\mathrm{max\_iter}$ do
3:     Update $\mathbf{W}_{\mathrm{r}}^{(n+1)}$ using (52)
4:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}$ using (53)
5:     Solve optimisation sub-problem (57) to update $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ using the projected gradient descent
6:     Project $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ onto feasible set
7:     Update $\mathbf{U}_{\mathrm{r}}^{(n+1)}$ using (55)
8:     Update $\mathbf{U}_{\mathrm{c}}^{(n+1)}$ using (56)
9:     if $\left\|\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right\|_{F}<\varepsilon$ and $\left\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right\|_{F}<\varepsilon$ then
10:        break
11:     end if
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{W}_{\mathrm{r}}^{*}=\mathbf{W}_{\mathrm{r}}^{(n)}$, $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
VIII Complexity Analysis
VIII-1 Projected Gradient Method with Penalty
The projected gradient method with adaptive penalties for radar waveform optimisation, defined in Algorithm 1, has an estimated per-iteration complexity as [42, 41],
$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),$ (65)
where $I$ is the number of iterations, $E$ denotes the cost of evaluating the objective function and its gradients, $n$ represents the number of optimisation variables in the tensor $\dot{\mathbf{W}}_{\mathrm{r}}$, $m$ is the number of constraints, and $k$ is the number of backtracking line search iterations per step. The dominant $n^{3}$ term arises from matrix operations in the projection step.
VIII-2 Gradient-assisted IPM
For communication waveform optimisation, we utilise a gradient-assisted Interior Point Method as outlined in Algorithm 2. Its per-iteration complexity is [44],
$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),$ (66)
where $n$ denotes the number of optimisation variables in the matrix $\mathbf{W}_{\mathrm{c}}$. Again, while the dominant $n^{3}$ term reflects matrix operations.
VIII-3 ADMM
The ADMM algorithm for unified optimisation defined in Algorithm 3, has per-iteration complexity as [45],
$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),$ (67)
where $n$ represents the combined number of optimisation variables, and $J$ is the number of inner projected gradient iterations per ADMM step. The dominant $n^{3}$ term arises from matrix operations in the projection steps.
IX Numerical Results
In this section, we present simulation results for our conventional benchmarks and the three previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed. A pathloss exponent of $\eta=3$ is considered to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is fixed at $\sigma_{h}^{2}=1$, and the probability of false alarm $P_{FA}$ is set to $10^{-4}$. $\overline{C}_{k}$ is set to $210^{-}$ 2 which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits.
IX-A Trade-off for the conventional benchmark
The radar subsystem employs a CIC matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced ZF.
Figure 1: The ISAC system performance at $P_{\mathrm{r}}=0.5$. a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
[…]
IX-B Radar Waveform KLD-based Optimisation
The radar waveform KLD-based optimisation algorithm is set with a maximum number of iterations of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
Figure 2: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using radar waveform optimisation: K-ROP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig. 2 illustrates the performance of both subsystems using the KLD-based radar waveform optimisation (K-ROP) technique. In Fig. 2.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ on the right axis and detection probability $P_{D}$ on the left axis versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance. By comparing the results here with the ZF-based design in Fig. 1, it can be noticed that there is a significant gain can be achieved using K-ROB. For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig. 1, demonstrating improved detection capabilities through K-ROP optimisation technique. For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system.
Fig. 2.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ on the right axis and BER on the left axis versus $P_{\mathrm{R}}/N_{o}$. For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB suggests it has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced. The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted from Fig. 1. These results highlight the intricate relationship between radar detection capabilities and communication system performance.
IX-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$. The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
Figure 3: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using communication waveform optimisation: K-COP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig. 3 illustrates the performance of both subsystems using the KLD-based communication waveform optimisation (K-COP) technique. In Fig. 3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ (right axis) and detection probability $P_{D}$ (left axis) versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform. At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig. 1. K-COP demonstrates improved detection capabilities over non-optimisation but remains below the radar waveform optimisation performance achieved using K-ROP in Fig. 2, which is $\{98,89.2,99.9\}\%$.
Fig. 3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ (right axis) and BER (left axis) versus $P_{\mathrm{R}}/N_{o}$. For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig. 1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs. K-COP optimisation technique provides significant performance gains over the non-optimised ZF scenario, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec. IX-B due to radar interference. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
IX-D Integrated Waveform KLD-based Optimisation for ISAC
The integrated waveform optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, $\rho=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
Figure 4: The ISAC system performance using the integrated waveform optimisation. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig. 4 illustrates the ISAC system performance under the KLD-based integrated waveform optimisation (K-IOP). In Fig. 4.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ (left axis) and $\mathrm{KLD}_{\mathrm{r}}$ (right axis) for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$. As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP. For instance, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, the achievable $\mathrm{KLD}_{\mathrm{r}}$ values for $\{\mathrm{T}_{1},\mathrm{T}_{2},\mathrm{T}_{3}\}$ are approximately $\{38.87,32.77,41.85\}$ bits, with corresponding detection probabilities $\{0.94,0.89,0.97\}$. Performance continues to rise at higher $\mathit{P}_{\mathrm{R}}/N_{o}$, achieving near-unity detection across all targets once $\mathit{P}_{\mathrm{R}}/N_{o}$ exceeds 22 dB.
In Fig. 4.b, the communication subsystem performance is depicted via the BER (left axis) and $\mathrm{KLD}_{\mathrm{c}}$ (right axis) for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$. Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER. For example, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, $\mathrm{KLD}_{\mathrm{c}}$ values of $\{77.19,28.13,260.53\}$ bits are observed for $\{\mathrm{UE}_{1},\mathrm{UE}_{2},\mathrm{UE}_{3}\}$, with corresponding BERs of approximately $\{5.1\times 10^{-9},1.7\times 10^{-4},\lesssim 10^{-10}\}$. Beyond 14-16 dB, the BER for all UEs drops rapidly– often reaching effectively error-free communication– and $\mathrm{KLD}_{\mathrm{c}}$ grows significantly as the interference from the radar subsystem is jointly minimised. Comparing these K-IOP results with the non-optimised ZF baseline (cf. Fig.1) or the separate (radar-only or communication-only) waveform approaches in Sec. IX-BIX-C, it is evident that the integrated design offers substantially improved performance simultaneously for radar detection and communication reliability. This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
Figure 5: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
Fig. 5 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for K-ROP, K-COP, and K-IOP optimisation techniques. At low $P_{\mathrm{R}}/N_{o}$ values (00-$10$ dB), K-IOP immediately achieves lower MSE than both K-ROP and K-COP. As $P_{\mathrm{R}}/N_{o}$ increases beyond 10 dB, the techniques begin to show more distinct performance. K-ROP demonstrates faster MSE decay, particularly in the $15$-$20$ dB range, while K-COP exhibits a steadier improvement but converges slightly more slowly. The K-IOP continues to outperform both separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values. Eventually, all three techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels. Notably, this MSE performance strongly correlates with the detection capabilities in Figs. 3.a, 2.a, and 4, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$. Conversely, missed detections or false alarms degrade the estimation by introducing noise-based or zero-valued entries.
Figure 6: Average CPU time (seconds) versus antenna array size $N$ for both optimisation techniques.
Fig. 6 presents the average CPU time versus antenna array size $N$, comparing all optimisation techniques. The K-ROP consistently demonstrates the lowest computational cost, with CPU times increasing from $0.001$ s at $N=10$ to $7.98$s at $N=100$. Meanwhile, K-COP exhibits moderate requirements, ranging from $0.04$ s to $14.07$s across the same $N$. By contrast, the K-IOP incurs the highest computational load, growing from $2.44$ s at $N=10$ to $37.69$s at $N=100$. For instance, at $N=50$, K-ROP requires only $2.80$s, whereas K-COP needs $3.90$s, and the K-IOP demands $19.96$s. These results indicate that while the radar-only and communication-only algorithms remain relatively efficient as $N$ increases, the K-IOP entails modestly higher computational complexity, while offering much higher performance.
X Conclusion
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP for ISAC approach, each demonstrated significant improvements over the non-optimised scenario. K-ROP yielded substantial enhancements in both target detection and communication performance while maintaining efficient computational requirements, as evidenced by the numerical results. K-COP, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels. Both techniques exhibited robust performance across varying SNR values. In addition, the K-IOP effectively unifies these advantages, consistently outperforming the two separate approaches in terms of detection probability, BER, and estimation accuracy. By jointly designing radar and communication waveforms, the K-IOP approach leverages the synergy between sensing and data transmission, thereby achieving higher overall KLD and a more balanced performance for future ISAC networks. Moreover, the numerical results revealed that exploiting the communication signal as part of the radar subsystem significantly boosts ISAC system capabilities, underscoring the importance of holistic design in shared-antenna deployments. These findings provide valuable insights for designing and optimising ISAC systems, offering a promising pathway to balance sensing and communication requirements in next-generation wireless networks. The proposed framework offers the flexibility to adapt system resources for diverse operating regimes, paving the way for more efficient and versatile wireless communications and sensing technologies.
Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimisation problem. The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.$ (68)
Applying the chain rule to (27), and substituting (68) we get,
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},$ (69)
where $\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{H}_{t}^{H}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$. Substituting this into (69) and simplifying, we obtain,
$\!\!\!\!\!\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!$ (70)
Finally, the gradient of $f$ with respect to $\mathbf{W}_{\mathrm{r},t}$, derived from (31) and (70), is,
$\!\!\!\!\!\![\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!$ (71)
For $\mathrm{KLD}_{\mathrm{c},k}$ gradient calculations, we consider how changes in $\mathbf{W}_{\mathrm{r}}$ affect $\sigma_{\eta}^{2}$, where it is defined as follows,
$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}$ (72)
The derivative of $\sigma_{\eta}^{2}$ with respect to $\mathbf{W}_{\mathrm{r},l}$ is,
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.$ (73)
Computing $\partial\mathrm{KLD}_{\mathrm{c},k}/\partial\sigma_{\eta}^{2}$,
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.$ (74)
Utilising the chain rule and (73), (74), we obtain,
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.$ (75)
Therefore, the gradient of $\mathrm{KLD}_{\mathrm{c},k}$ with respect to $\mathbf{\dot{W}}_{\mathrm{r}}$ is,
$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (76)
Equations (71) and (76) provide the mathematical foundation for the gradients used in our optimisation algorithm.